# Foliations by Critical Surfaces of the Hawking Energy in Asymptotically Flat Initial Data Sets

### v1

@@ 2.1. Foliation construction
Note that $W_{2}(\Sigma_{\xi,r})=\mathcal{O}(r^{-5})$, then the proof follows roughly as in the case $k=0$ of [9, Lemma 20].
We have by the decay of the metric
@@
$W_{1}(\Sigma_{\xi,r})+W_{2}(\Sigma_{\xi,r})+\lambda H(\Sigma_{\xi,r})=Y_{1}\in\Lambda_{1}$
Thus,
@@
Then as in [9, Lemma 21], it is direct to see that:
@@
With this result, finding critical surfaces of the Hawking energy that are deformations of coordinate spheres reduces to identifying critical points of the function $G_{r}$. To locate such critical points, we require a more explicit form of $G_{r}$. In [9, Lemma 22], the asymptotic expansion of $G_{r}$ was computed in the absence of the tensor $k$. In our case, a similar expansion is derived using the same procedure, but with an additional term accounting for $k$.
@@
Now we can obtain the following result which is analogous to [9, Lemma 24].
@@
There exist $\tau>0$, $\tilde{\delta}_{0}\in(0,1/2)$, and $r_{0}>1$, $\tilde{\eta}_{i}>0$ for $i=1,2,3$ depending only on $g$ and $k$ such that, provided $r>r_{0}$ and $\tilde{\eta}_{i}>\eta_{i}$,
@@
By using the same argument used in [9, Lemma 24] to estimate $\bar{D}_{\xi}G_{r,2}$ we obtain
@@
with $\alpha=r\delta(\frac{\xi}{|\xi|},\nu)$ and where since $\pi=k-\operatorname{tr}kg$, $P=-\pi(\nu,\nu)$. Now by using a Taylor expansion around the origin, we have
@@
Where $\tilde{\xi}$ is a constant vector such that $0\leq|\tilde{\xi}|\leq|\xi|$, and by (12), the second term of the integral can be written as
(25) $r^{2}\int_{S_{r}(0)}2DW_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),r\delta(\xi,\nu)\right)d\mu_{\delta}$
@@
(26) $\begin{split}-r^{2}\int_{S_{r}(0)}2W_{2}\alpha d\mu_{\delta}&=-\mathcal{O}(\eta_{3}^{2})\\ -r^{2}\int_{S_{r}(0)}2DW_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),r\delta(\xi,\nu)\right)d\mu_{\delta}&=-\mathcal{O}(\eta_{3}|\xi|)\\ -\frac{r^{2}}{2}\bar{D}^{2}\left(\int_{S_{r}(\xi r)}2W_{2}\alpha d\mu_{\delta}\right)_{|\xi=\tilde{\xi}}(r\xi,r\xi)&=-\mathcal{O}(\eta_{3}|\xi|^{2}).\end{split}$
@@
For convexity, as in [9, Lemma 24] one sees that
@@
We follow the structure of the proof in [9]. First, we show that for every $r>1$ sufficiently large, there exists a stable area-constrained critical sphere with area radius $r$. To see this, we decompose
@@
Let $a\in\mathbb{R}^{3}$. Differentiating the identity $(D\overline{G}_{r})|_{\xi(r)}(a)=0$ with respect to $r$, we obtain
(37) $(\overline{D}^{2}G_{r})|_{\xi(r)}(a,\xi^{\prime}(r))+(\overline{D}G^{\prime}_{r})|_{\xi(r)}(a)=0.$
We need to estimate $(\overline{D}G^{\prime}_{r})|_{\xi(r)}(a)$, then proceeding as in [9, Proposition 49] one finds
(38) $(\overline{D}G^{\prime}_{2,r})|_{\xi(r)}(a)=\mathcal{O}((\eta_{1}+\eta_{2})r^{-1})$
@@
$\tau|\xi^{\prime}(r)|^{2}\leq(\overline{D}^{2}G_{r})|_{\xi(r)}(\xi^{\prime}(r),\xi^{\prime}(r))\leq-(\overline{D}G^{\prime}_{r})|_{\xi(r)}(\xi^{\prime})=\mathcal{O}((\eta_{1}+\eta_{2}+\eta_{3})|\xi^{\prime}(r)|r^{-1}).$
@@
$\frac{\partial\Phi}{\partial r}=y+\xi(r)+r\xi^{\prime}(r)+\overline{D}_{\xi^{\prime}(r)}u_{\xi,r}y+u_{\xi,r}^{\prime}(y)y$
@@
(41) $\begin{split}\delta(\frac{\partial\Phi}{\partial r},y)&=1+\delta(\xi(r),y)+r\delta(\xi^{\prime}(r),y)+\overline{D}_{\xi^{\prime}(r)}u_{\xi,r}+u_{\xi,r}^{\prime}(y)\\ &\geq 1-|\xi(r)|+\mathcal{O}(\eta_{1}+\eta_{2}+\eta_{3})+\mathcal{O}(r^{-1})\end{split}$
@@
Note that compared to [9] we relaxed the assumptions on $\mathrm{Sc}$ by changing the little $o$ bounds by big $\mathcal{O}$ bound with a small constant $\eta_{i}$. By relaxing the assumption this way we obtain that $\lim_{r\to\infty}\xi(r)$ is not necessarily zero, meaning that the surfaces are not centered and their center may slightly move while preserving the foliation property. Also, the main difference with the assumptions for the foliation constructed in [10] is that there the technique of Lamm, Metzger and Schulze [20] was used. This requires stronger conditions on the decay of the metric and on the scalar curvature for which a decay of the form $\mathrm{Sc}=\mathcal{O}(\eta|x|^{-5})$ for small $\eta$ was required.
We also obtain a uniqueness result with a proof analogous to the one of [9, Theorem 8] which we include for completeness.
@@
and $\Sigma_{j}\neq\Sigma(r_{j})$ where $r_{j}=r(\Sigma_{j})$. For $\epsilon_{0}>0$ small enough (and as $k$ decays to zero), Lemma 4.2 in [18] implies the uniform, scale-invariant curvature estimate
$\|B\|_{L^{\infty}(\Sigma_{j})}=\mathcal{O}(r_{j}^{-1})$
with corresponding higher-order estimates as well as the estimate
$\lambda(\Sigma_{j})=\mathcal{O}(r_{j}^{-3}).$
Passing to a subsequence, if necessary, the rescaled surfaces $\widetilde{\Sigma}_{j}=r_{j}^{-1}\Sigma_{j}$ converge smoothly to a Euclidean Willmore surface $\widetilde{\Sigma}\subset\mathbb{R}^{3}$ satisfying
$\int_{\widetilde{\Sigma}}|\mathring{B}_{\delta}|^{2}\,d\mu_{\delta}<\epsilon_{0},\qquad|\widetilde{\Sigma}|=4\pi,\qquad\frac{1}{4\pi}\int_{\widetilde{\Sigma}}y\,d\mu_{\delta}=\xi_{0}$
By [19, Theorem 2.7] $\widetilde{\Sigma}=S_{1}(\xi_{0})$. It follows that $\Sigma_{j}$ is a perturbation of a coordinate sphere for large $j$. Then for $j$ large, $W_{1}(\Sigma_{j})+W_{2}(\Sigma_{j})+\lambda H(\Sigma_{j})=0\in\Lambda_{1}(S_{\xi,r_{j}})$ for certain $\xi$ and $|\Sigma_{j}|=4\pi r_{j}^{2}$, then by the uniqueness statement of the Lyapunov-Schmidt reduction of Proposition 2.3, $\Sigma_{j}$ is one of the surfaces obtained there in the sense that $\Sigma_{j}=\Sigma_{\xi_{j},r_{j}}$ for some $\xi_{j}$, but as $\Sigma_{\xi_{j},r_{j}}$ is an area-constrained critical surface, then $\xi_{j}$ is also the only critical point of $G_{r_{j}}$. This implies that $\xi_{j}=\xi(r_{j})$ getting a contradiction. ∎
@@
Note that compared to the case in [9] $\xi_{0}$ is not equal to zero, but this does not compromise the uniqueness result since it only relies on the surfaces being a perturbation of coordinate spheres so that one can use the uniqueness of the Lyapunov-Schmidt reduction.
We can relax the conditions on $\mathrm{Sc}$ and $k$, obtaining a family of surfaces at infinity, similar to [8, Theorem 6], but not a foliation. This means that the surfaces may intersect with each other.
@@
(42) $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4})\quad\text{and}\quad\pi(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-2})$
where $\pi=k-\operatorname{tr}k\,g$ and $\rho=\frac{x^{i}}{|x|}\partial_{i}$. Then there exists a compact set $K\subset M$, a number $\lambda_{0}>0$, and a family $\{\Sigma(\lambda):\lambda\in(0,\lambda_{0})\}$ of on-center stable area-constrained critical spheres $\Sigma(\lambda)$ of the Hawking functional, satisfying (4) with parameter $\lambda$ such that as $\lambda\to 0$, $|\Sigma(\lambda)|\to\infty$ and $R(\lambda)\to\infty$, where $R(\lambda):=\sup_{x\in\Sigma(\lambda)}|x|$.
@@
(43) $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2}),\quad k(\cdot,\rho)=\mathcal{O}(\eta_{2}|x|^{-2})\quad\text{and}\quad(\nabla_{\rho}\pi)(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-3}),$
then the family of surfaces is unique in the sense that there exist a small constant $\epsilon_{0}>0$ and a compact set $K\subset M$ which depend only on $(M,g,k)$ such that for every $\tilde{\delta}>0$, there exists a large constant $r_{0}>1$ such that every area-constrained critical sphere $\Sigma\subset M\setminus K$ of the Hawking functional with
@@
$G_{r,2}(\xi)=2r\int_{|x|>r|\xi|}\mathrm{Sc}(x)\,dv_{\delta},\qquad G_{r,3}(\xi)=-r^{2}\int_{S_{r}(r\xi)}\pi(\nu,\nu)^{2}\,d\mu_{\delta},$
By $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4})$
@@
and by using a Taylor expansion for the center of the integrals
(45) $\begin{split}-r^{2}\int_{S_{r}(r\xi)}\pi(\nu,\nu)^{2}d\mu_{\delta}=&-r^{2}\int_{S_{r}(0)}\pi(\nu,\nu)^{2}d\mu_{\delta}-r^{2}\int_{S_{r}(0)}2W_{2}\,\delta(r\xi,\nu)d\mu_{\delta}\\ &-r^{2}\int_{S_{r}(r\tilde{\xi})}DW_{2}\left(r\delta(\xi,\nu),r\delta(\xi,\nu)\right)d\mu_{\delta}\\ &\geq-\mathcal{O}(\eta_{2}^{2})-\mathcal{O}(\eta_{2}|\xi|)-\mathcal{O}(|\xi|^{2})\end{split}$
where $\tilde{\xi}$ is such that $0\leq|\tilde{\xi}|\leq|\xi|$ and we used that on $S_{r}(0)$, $\nu=\rho$. We know that $G_{r}(\xi)$ is continuous and that $G_{1}$ is strictly convex and blows up as $|\xi|\nearrow 1$. These bounds imply that for large $r$ the continuous function $G_{r}$ attains a global minimum on the compact set $\{|\xi|\leq 1-\tilde{\delta}\}$ and the minimizer cannot occur at the boundary $|\xi|=1-\tilde{\delta}$ (for $|\xi|$ near 1 the blow-up of $G_{1}$ dominates the small perturbations). Hence for every large $r$ there exists at least one interior minimizer $\xi(r)$ and the surface $\Sigma_{\xi(r),r}$ is an area-constrained critical sphere with area $4\pi r^{2}$. Note that this minimum is not necessarily unique, to guarantee uniqueness, we need convexity of $G_{r}(\xi)$.
Now assuming (43), we will show that $D^{2}G_{r}$ is positive definite. By [8, Lemma 20, Corollary 21] we have that $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2})$ implies that $D^{2}G_{r,2}$ is positive semidefinite up to an error $\mathcal{O}(\eta_{2})$, i.e. $D^{2}G_{r,2}\geq-c\eta_{2}\text{Id}$. For $D^{2}G_{r,3}$ we use again an expansion at the center of the integrals obtaining
(46) $\begin{split}D^{2}G_{r,3}(a,a)=&-r^{2}\int_{S_{r}(r\xi)}2DW_{2}\left(r\delta(a,\nu),r\delta(a,\nu)\right)d\mu_{\delta}\\ =&-r^{2}\int_{S_{r}(0)}2DW_{2}\left(r\delta(a,\nu),r\delta(a,\nu)\right)d\mu_{\delta}+\mathcal{O}(|\xi|)\\ =&\mathcal{O}(\eta_{2})+\mathcal{O}(|\xi|),\end{split}$
where we used that each term of the second variation depends on the radial directions, $\pi(\rho,\rho)$ $(\nabla_{\rho}\pi)(\rho,\rho)$ or $k(\rho,\cdot)$. Then we have that for $\eta_{i}$ and $|\xi|$ small enough $D^{2}G_{r}$ is positive definite. Repeating the argument used in Theorem 2.12 we obtain the uniqueness. ∎
@@
Recall that $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4})$ is implied by $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2})$ as shown in (34). The reason why we cannot guarantee a foliation in the previous result is that, with our assumptions, we don’t have good control over $|\xi|$. When going through the proof of Lemma 2.9, and trying to get the estimate for $|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,3})(\xi)$, note that all the terms in (26) except the last one depend on the radial directions, $\pi(\rho,\rho)$ $(\nabla_{\rho}\pi)(\rho,\rho)$ and $k(\rho,\cdot)$. With this, we can see that $|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,3})(\xi)\geq-\mathcal{O}(\eta_{2}^{2})-\mathcal{O}(\eta_{2}|\xi|)-\mathcal{O}(|\xi|^{2})$. Similarly we have $|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,2})(\xi)\geq-\mathcal{O}(\eta_{1})-\mathcal{O}(\eta_{1}|\xi|)-\mathcal{O}(|\xi|^{2})$. Then we have a worst estimate on $\bar{D}_{\xi}G_{r}$ than (28):
@@
and because of (47), we do not have an estimate on $|\xi|$ like in (36). Therefore we have a worst control on $(\overline{D}G^{\prime}_{3,r})|_{\xi(r)}(a)$ (see (39)), which also implies that we don’t have a good estimate on $|\xi^{\prime}|$ which was crucial to show that the family of surfaces form a foliation.

### v2

@@ 2.1. Foliation construction
Note that $W_{2}(\Sigma_{\xi,r})=\mathcal{O}(r^{-5})$, then the proof follows roughly as in the case $k=0$ of [eichko, Lemma 20].
We denote by $^{\delta}\Delta_{\xi,r}$ the laplacian on $S_{\xi,r}$ with respect to the Euclidean metric $\delta$. We have by the decay of the metric
@@
$W_{1}(\Sigma_{\xi,r})+W_{2}(\Sigma_{\xi,r})+\lambda H(\Sigma_{\xi,r})=Y_{1}\in\Lambda_{1}.$
Then, using that by the decay of $k$, $W_{2}(\Sigma_{\xi,r})=\mathcal{O}(r^{-5})$ and
@@
Then as in [eichko, Lemma 21], it is direct to see that:
@@
With this result, finding critical surfaces of the Hawking energy that are deformations of coordinate spheres reduces to identifying critical points of the function $G_{r}$. To locate such critical points, we require a more explicit form of $G_{r}$. In [eichko, Lemma 22], the asymptotic expansion of $G_{r}$ was computed in the absence of the tensor $k$. In our case, a similar expansion is derived using the same procedure, but with an additional term accounting for $k$.
@@
Now we can obtain the following result, which is analogous to [eichko, Lemma 24].
@@
There exist constants $\tilde{\eta}_{i}>0$, i=1,2,3 such that if for $\eta_{i}<\tilde{\eta}_{i}$ it holds
and
then there exist $\tau>0$, $\tilde{\delta}_{0}\in(0,1/2)$, and $r_{0}>1$, $\tilde{\eta}_{i}>0$ for $i=1,2,3$ depending only on $g$ and $k$ such that, provided $r>r_{0}$ and $\tilde{\eta}_{i}>\eta_{i}$,
@@
By using the same argument used in [eichko, Lemma 24] to estimate $\bar{D}_{\xi}G_{r,2}$ we obtain
@@
with $\alpha=r\delta(\frac{\xi}{|\xi|},\nu)$ and where since $\pi=k-\operatorname{tr}kg$, $P=-\pi(\nu,\nu)$.
Now we consider a Taylor expansion around the origin of the integral expression, that is, we consider the function $f(\xi)=-r^{2}\int_{S_{r}(\xi r)}2W_{2}\alpha d\mu_{\delta}$ Then, applying Taylor’s theorem at $\xi=0$, we obtain
@@
where $\tilde{\xi}$ is a point on the segment between 0 and $\xi$ (i.e. $\tilde{\xi}=t\xi$ for some $t\in(0,1)$), and by (16), the second term of the integral can be written as
(31) $r^{2}\int_{S_{r}(0)}2\mathcal{D}W_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),r\delta(\xi,\nu)\right)d\mu_{\delta}.$
@@
(32) $\begin{split}-r^{2}\int_{S_{r}(0)}2W_{2}\alpha d\mu_{\delta}&=-\mathcal{O}(\eta_{3}^{2})\\ -r^{2}\int_{S_{r}(0)}2\mathcal{D}W_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),r\delta(\xi,\nu)\right)d\mu_{\delta}&=-\mathcal{O}(\eta_{3}|\xi|)\\ -\frac{r^{2}}{2}\bar{D}^{2}\left(\int_{S_{r}(\xi r)}2W_{2}\alpha d\mu_{\delta}\right)_{|\xi=\tilde{\xi}}(r\xi,r\xi)&=-\mathcal{O}(\eta_{3}|\xi|^{2}).\end{split}$
@@
For convexity, as in [eichko, Lemma 24] one sees that
@@
We follow the structure of the proof in [eichko]. First, we show that for every $r>1$ sufficiently large, there exists a stable area-constrained critical sphere with area radius $r$. To see this, we decompose
@@
Let $a\in\mathbb{R}^{3}$. Differentiating the identity $(\bar{D}G_{r})|_{\xi(r)}(a)=0$ with respect to $r$, we obtain
(43) $(\bar{D}^{2}G_{r})|_{\xi(r)}(a,\xi^{\prime}(r))+(\bar{D}G^{\prime}_{r})|_{\xi(r)}(a)=0.$
We need to estimate $(\bar{D}G^{\prime}_{r})|_{\xi(r)}(a)$, then proceeding as in [eichko, Proposition 49] one finds
(44) $(\bar{D}G^{\prime}_{2,r})|_{\xi(r)}(a)=\mathcal{O}((\eta_{1}+\eta_{2})r^{-1})$
@@
$\tau|\xi^{\prime}(r)|^{2}\leq(\bar{D}^{2}G_{r})|_{\xi(r)}(\xi^{\prime}(r),\xi^{\prime}(r))\leq-(\bar{D}G^{\prime}_{r})|_{\xi(r)}(\xi^{\prime})=\mathcal{O}((\eta_{1}+\eta_{2}+\eta_{3})|\xi^{\prime}(r)|r^{-1}).$
@@
$\frac{\partial\Phi}{\partial r}=y+\xi(r)+r\xi^{\prime}(r)+\bar{D}_{\xi^{\prime}(r)}u_{\xi,r}y+u_{\xi,r}^{\prime}(y)y$
@@
(47) $\begin{split}\delta(\frac{\partial\Phi}{\partial r},y)&=1+\delta(\xi(r),y)+r\delta(\xi^{\prime}(r),y)+\bar{D}_{\xi^{\prime}(r)}u_{\xi,r}+u_{\xi,r}^{\prime}(y)\\ &\geq 1-|\xi(r)|+\mathcal{O}(\eta_{1}+\eta_{2}+\eta_{3})+\mathcal{O}(r^{-1})\end{split}$
@@
Note that compared to [eichko] we relaxed the assumptions on $\mathrm{Sc}$ by changing the little $o$ bounds by big $\mathcal{O}$ bound with a small constant $\eta_{i}$. By relaxing the assumption this way we obtain that $\lim_{r\to\infty}\xi(r)$ is not necessarily zero, meaning that the surfaces are not centered and their center may slightly move while preserving the foliation property. Also, the main difference with the assumptions for the foliation constructed in [Friedrich2020] is that there the technique of Lamm, Metzger and Schulze [willflat] was used. This requires stronger conditions on the decay of the metric and on the scalar curvature for which a decay of the form $\mathrm{Sc}=\mathcal{O}(\eta|x|^{-5})$ for small $\eta$ was required.
We also obtain a uniqueness result with a proof analogous to the one of [eichko, Theorem 8] which we include for completeness.
@@
The argument follows closely the proof of [eichko, Theorem 8], with the curvature control replaced by the scale-invariant estimates proved in Eichmair–Koerber–Metzger–Schulze [eichmair2024huisken].
@@
and $\Sigma_{j}\neq\Sigma(r_{j})$ where $r_{j}=r(\Sigma_{j})$.
Since the tensor $k$ decays as $\mathcal{O}(|x|^{-2})$ and its derivatives as $\mathcal{O}(|x|^{-3})$, the Hawking operator differs from the Willmore operator by lower–order perturbations. Hence the scale–invariant curvature bounds for large area–constrained Willmore spheres from [eichmair2024huisken, Lemmas 6 and 8, Proposition 10, Corollary 14] apply to the present situation with only lower–order error terms. In particular, for each sufficiently large $j$,
(48) $|x|\big|B-r_{j}^{-1}g_{|\Sigma_{j}}\big|+|x|^{2}\,|\nabla B|\;=\;\mathcal{O}(r_{j}^{-1/2}+\frac{1}{\min_{x\in\Sigma_{j}}|x|}),$
By our assumptions $|x|\simeq r_{j}$ on $\Sigma_{j}$. Substituting these comparabilities into (48) yields
(49) $\|B\|_{L^{\infty}(\Sigma_{j})}=\mathcal{O}(r_{j}^{-1}),\qquad\|\nabla B\|_{L^{\infty}(\Sigma_{j})}=\mathcal{O}(r_{j}^{-2}),$
We can also get an estimate on the Lagrange parameter $\lambda(\Sigma_{j})$. Integrating the Euler–Lagrange equation (4) over $\Sigma_{j}$ and using $\int_{\Sigma_{j}}\Delta_{\Sigma_{j}}H\,d\mu=0$ and $\int_{\Sigma_{j}}\mathrm{div}_{\Sigma_{j}}X\,d\mu=0$, we obtain
$\lambda(\Sigma_{j})\int_{\Sigma_{j}}H\,d\mu=-\!\int_{\Sigma_{j}}\!\Big(H|\mathring{B}|^{2}+H\,\mathrm{Ric}(\nu,\nu)+P(\nabla_{\nu}\operatorname{tr}k-\nabla_{\nu}k(\nu,\nu))+\tfrac{1}{2}HP^{2}-2\,k(\nabla^{\Sigma}P,\nu)\Big)\,d\mu.$
Using (49), $H=2r_{j}^{-1}+\mathcal{O}(r_{j}^{-2})$, $|P|=\mathcal{O}(r_{j}^{-2})$, $|\nabla P|=\mathcal{O}(r_{j}^{-3})$, $\mathrm{Ric}=\mathcal{O}(r_{j}^{-3})$, and $|\Sigma_{j}|=4\pi r_{j}^{2}+\mathcal{O}(1)$, we can estimate the right hand side integral by $\mathcal{O}(r_{j}^{-3})$. Since $\int_{\Sigma_{j}}H\,d\mu=8\pi r_{j}+\mathcal{O}(1)\simeq c\,r_{j}$, it follows that
(50) $\lambda(\Sigma_{j})=\mathcal{O}(r_{j}^{-3}).$
Define the rescaled surfaces $\tilde{\Sigma}_{j}:=r_{j}^{-1}\Sigma_{j}\subset(\mathbb{R}^{3},\delta)$. By (49), the family $\{\tilde{\Sigma}_{j}\}$ has uniformly bounded curvature and area, and hence (after passing to a subsequence) converges smoothly to a closed Willmore surface $\tilde{\Sigma}\subset\mathbb{R}^{3}$ satisfying
$\int_{\tilde{\Sigma}}|\mathring{B}_{\delta}|^{2}\,d\mu_{\delta}<\varepsilon_{0},\qquad|\tilde{\Sigma}|=4\pi,\qquad\frac{1}{4\pi}\int_{\tilde{\Sigma}}y\,d\mu_{\delta}=\xi_{0}.$
By [kuwert2001willmore, Theorem 2.7] (or also by the quantitative rigidity results of De Lellis–Müller [de2005optimal, Theorem 1] and [de2006ac, Theorem 2]), it follows that $\tilde{\Sigma}=S_{1}(\xi_{0})$ is a round sphere.
For $j$ large, each $\Sigma_{j}$ is therefore a small normal graph over a coordinate sphere:
$\Sigma_{j}=\Sigma_{\xi_{j},r_{j}}(u_{j})\quad\text{with}\quad\|u_{j}\|_{C^{4}(S^{2})}\leq\varepsilon,$
for some small $\varepsilon>0$ depending only on $\varepsilon_{0}$. Then for $j$ large, $W_{1}(\Sigma_{j})+W_{2}(\Sigma_{j})+\lambda H(\Sigma_{j})=0\in\Lambda_{1}(S_{\xi,r_{j}})$ for certain $\xi$ and $|\Sigma_{j}|=4\pi r_{j}^{2}$, then by the uniqueness statement of the Lyapunov-Schmidt reduction of Proposition 2.3, $\Sigma_{j}$ is one of the surfaces obtained there in the sense that $\Sigma_{j}=\Sigma_{\xi_{j},r_{j}}$ for some $\xi_{j}$, but as $\Sigma_{\xi_{j},r_{j}}$ is an area-constrained critical surface, then $\xi_{j}$ is also the only critical point of $G_{r_{j}}$. This implies that $\xi_{j}=\xi(r_{j})$ getting a contradiction. ∎
@@
Note that compared to the case in [eichko] $\xi_{0}$ is not equal to zero, but this does not compromise the uniqueness result since it only relies on the surfaces being a perturbation of coordinate spheres so that one can use the uniqueness of the Lyapunov-Schmidt reduction.
We can relax the conditions on $\mathrm{Sc}$ and $k$, obtaining a family of surfaces at infinity, similar to [willcen, Theorem 6], but not a foliation. This means that the surfaces may intersect with each other.
@@
(51) $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4}).$
Then there exists a compact set $K\subset M$, a number $\lambda_{0}>0$, and a family $\{\Sigma(\lambda):\lambda\in(0,\lambda_{0})\}$ of on-center stable area-constrained critical spheres $\Sigma(\lambda)$ of the Hawking functional, satisfying (4) with parameter $\lambda$ such that as $\lambda\to 0$, $|\Sigma(\lambda)|\to\infty$ and $R(\lambda)\to\infty$, where $R(\lambda):=\sup_{x\in\Sigma(\lambda)}|x|$.
@@
(52) $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2}),$
(53) $\pi(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-2}),\quad k(\cdot,\rho)=\mathcal{O}(\eta_{2}|x|^{-2})\quad\text{and}\quad(\nabla_{\rho}\pi)(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-3}),$
where $\pi=k-\operatorname{tr}k\,g$ and $\rho=\frac{x^{i}}{|x|}\partial_{i}$. Then the family of surfaces is unique in the sense that there exist a small constant $\epsilon_{0}>0$ and a compact set $K\subset M$ which depend only on $(M,g,k)$ such that for every $\tilde{\delta}>0$, there exists a large constant $r_{0}>1$ such that every area-constrained critical sphere $\Sigma\subset M\setminus K$ of the Hawking functional with
@@
$G_{r,2}(\xi)=2r\int_{|x|>r|\xi|}\mathrm{Sc}(x)\,dv_{\delta},\qquad G_{r,3}(\xi)=-r^{2}\int_{S_{r}(r\xi)}\pi(\nu,\nu)^{2}\,d\mu_{\delta}.$
By the decay of $k$ it holds $G_{r,3}(\xi)=\mathcal{O}(1)$ and by $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4})$
@@
We know that $G_{r}(\xi)$ is continuous and that $G_{1}$ is strictly convex and blows up as $|\xi|\nearrow 1$. The bounds for $G_{r,2}$ and $G_{r,3}$ imply that for large $r$ the continuous function $G_{r}$ attains a global minimum on the compact set $\{|\xi|\leq 1-\tilde{\delta}\}$ and the minimizer cannot occur at the boundary $|\xi|=1-\tilde{\delta}$ (for $|\xi|$ near 1 the blow-up of $G_{1}$ dominates the small perturbations). Hence for every large $r$ there exists at least one interior minimizer $\xi(r)$ and the surface $\Sigma_{\xi(r),r}$ is an area-constrained critical sphere with area $4\pi r^{2}$. Note that this minimum is not necessarily unique, to guarantee uniqueness, we need convexity of $G_{r}(\xi)$.
Now assuming (52, 53), we will show that $\bar{D}^{2}G_{r}$ is positive definite. By [willcen, Lemma 20, Corollary 21] we have that $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2})$ implies that $\bar{D}^{2}G_{r,2}$ is positive semidefinite up to an error $\mathcal{O}(\eta_{2})$, i.e. $\bar{D}^{2}G_{r,2}\geq-c\eta_{2}\text{Id}$.
For $\bar{D}^{2}G_{r,3}$, like we did for (30) we use again a Taylor expansion at $\xi=0$ obtaining
(55) $\begin{split}\bar{D}^{2}G_{r,3}(a,a)=&-r^{2}\int_{S_{r}(r\xi)}2\mathcal{D}W_{2}\left(r\delta(a,\nu),r\delta(a,\nu)\right)d\mu_{\delta}\\ =&-r^{2}\int_{S_{r}(0)}2\mathcal{D}W_{2}\left(r\delta(a,\nu),r\delta(a,\nu)\right)d\mu_{\delta}+\mathcal{O}(|\xi|)\\ =&\mathcal{O}(\eta_{2})+\mathcal{O}(|\xi|),\end{split}$
where we used that each term of the second variation depends on the radial directions, $\pi(\rho,\rho)$ $(\nabla_{\rho}\pi)(\rho,\rho)$ or $k(\rho,\cdot)$. Then we have that for $\eta_{i}$ and $|\xi|$ small enough $\bar{D}^{2}G_{r}$ is positive definite. Repeating the argument used in Theorem 2.12 we obtain the uniqueness. ∎
@@
$i)$ Recall that the curvature condition $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4})$ is implied by $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2}),$ as shown in (40).
ii)ii) Note that for this family of surfaces not only the outer radius diverges but also the inner radius $\inf_{x\in\Sigma(\lambda)}|x|$ does.
$iii)$ The key reason why, under these weaker assumptions, we cannot guarantee that the family of surfaces constructed above forms a foliation is the lack of control on the size of the translation parameter $|\xi|$.
Indeed, revisiting the proof of Lemma 2.9, the crucial estimate for the radial derivative $|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,3})(\xi)$ involves, via the expansion (32), the terms depending on the radial components of $k$ and $\pi$, namely $\pi(\rho,\rho)$, $(\nabla_{\rho}\pi)(\rho,\rho)$, and $k(\rho,\cdot)$. From this, one obtains
$|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,3})(\xi)\geq-\mathcal{O}(\eta_{2}^{2})-\mathcal{O}(\eta_{2}|\xi|)-\mathcal{O}(|\xi|^{2}),$
and similarly,
$|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,2})(\xi)\geq-\mathcal{O}(\eta_{1})-\mathcal{O}(\eta_{1}|\xi|)-\mathcal{O}(|\xi|^{2}).$
Combining these inequalities gives a weaker estimate for the total derivative:
@@
Compared with (34), this bound yields less precise control on the growth of $G_{r}$ with respect to $\xi$. In particular, we can no longer obtain an estimate for $|\xi|$ as in (42). This, in turn, results in a weaker bound on $(\bar{D}G^{\prime}_{3,r})|_{\xi(r)}(a)$ (see (45)), which also implies that we do not have a good estimate on $|\xi^{\prime}|$, a quantity that was crucial in proving that the family of surfaces forms a foliation.